find the mean of the 15 th through 55 th terms inclusive of the sequence:
a billow n = 7 - 4n please i need help to solve any one please help.....

Answer by RAY100(1637)

a(n) = 7-4n
.
The form is that of an arithmetic series, a(n) = a(1) +(n-1)d
.
a(15) = 7-4(15) = -53
.
a(55) = 7-4(55) = -213
.
Mean is average in arithmetic series
.
mean = - (213 +53) /2 = -133